# Contextualize-then-Aggregate: Circuits for In-Context Learning in Gemma-2 2B

### v1

@@ Abstract
In-Context Learning (ICL) is an intriguing ability of large language models (LLMs). Despite a substantial amount of work on its behavioral aspects and how it emerges in miniature setups, it remains unclear which mechanism assembles task information from the individual examples in a fewshot prompt. We use causal interventions to identify information flow in Gemma-2 2B for five naturalistic ICL tasks. We find that the model infers task information using a two-step strategy we call contextualize-then-aggregate: In the lower layers, the model builds up representations of individual fewshot examples, which are contextualized by preceding examples through connections between fewshot input and output tokens across the sequence. In the higher layers, these representations are aggregated to identify the task and prepare prediction of the next output. The importance of the contextualization step differs between tasks, and it may become more important in the presence of ambiguous examples. Overall, by providing rigorous causal analysis, our results shed light on the mechanisms through which ICL happens in language models.
@@ 1 Introduction
In-Context Learning (ICL) is an intriguing property of large language models and has spurred a substantial amount of interest into how transformers are able to perform it (e.g. Brown et al., 2020; Min et al., 2022a; Garg et al., 2022; Akyürek et al., 2023; Cho et al., 2025; Wang et al., 2023). Prior work in mechanistic interpretability has found function vectors (Todd et al., 2024; Hendel et al., 2023): attention heads whose output encodes task information, and which are causally responsible for the prediction of the response. However, the circuit by which these vectors are assembled remains only partly understood. Recent work has proposed that each fewshot example’s output token computes task information, which is then aggregated by attention heads to predict the next output (Wang et al., 2023; Cho et al., 2025; Kharlapenko et al., 2025), but this strategy has also been observed to leave part of the models’ performance unexplained (Cho et al., 2025).
In this paper, we use causal interventions (e.g. Vig et al., 2020; Geiger et al., 2021; Meng et al., 2022) to identify a circuit performing ICL on five naturalistic tasks (Capitalization, Country-Capital, Present-Past, Person-Sport, Copying) in Gemma-2 2B (Team et al., 2024). We replicate the relevance of the aggregation step suggested by prior work, but show that it explains only a portion of the model’s full performance. We use causal interventions to identify information flow that recovers at least 90% of the model’s performance. A key idea is to first identify a computation graph between the tokens in a prompt (Figure 1). We find that the required graph differs between tasks: whereas the simplest circuit is largely sufficient on some tasks, others require additional computation paths. We find that the representations of fewshot examples later in the prompt are contextualized by information from prior fewshot examples, particularly, the last preceding fewshot example. Using causal interventions, we show that contextualization transports information about the input and output spaces, and the task itself.
@@
    Obtain causally faithful information flow from the fewshot examples to the next-token prediction, explaining $\geq 90\%$ of the model’s performance.
    A key idea here is to focus on position-level circuits, where information flows only between a restricted subset of positions in a prompt, which is key to making the resulting computation graph interpretable. This is distinct from Kharlapenko et al. (2025) in providing information flow at the level of individual fewshot examples, rather than aggregating across them. It is distinct from Cho et al. (2025) in providing a computation graph supported by causal effects rather than representation similarity, recovering a large fraction of the model’s performance.
@@
Figure 1: An $N$-shot ICL prompt (here, $N=2$), at the example of the Country-Capital task. The green edges describe a Parallel subcircuit that assembles information at each fewshot example in parallel ($x_{i}\rightarrow y_{i}$) and then aggregates these at $t_{N+1}$. The blue edges describe a Contextualization subcircuit that contextualizes the representation of each fewshot. A circuit involving both components recovers most of Gemma2-2B’s performance; aggregating without contextualization leads to a breakdown in performance on this task. See Figure 24 for other tasks.
@@ 2 Background: Aggregation and Function Vectors
Todd et al. (2024) (followed by Yin and Steinhardt (2025)) operationalize Function Vector Heads as heads whose activations, when patched from valid ICL prompts to shuffled uninformative prompts, can most increase the likelihood of the desired target (i.e., better performing the task articulated by the valid prompts). Aggregating activations from these attention heads results in function vectors, which can trigger execution of the task in uninformative or zero-shot contexts. Relatedly, Kharlapenko et al. (2025) find sparse autoencoder (SAE) features that encode task information (in fewshot examples) and features that execute them (at $t_{N+1}$). However, it remains largely open what circuit assembles this information, and how information is assembled from the different fewshot examples to these function vectors. One simple hypothesis could be that each fewshot example computes a representation, and this is aggregated by Function Vector Heads.
On the other hand, function vectors might not play an important role in certain kind of ICL tasks. Cho et al. (2025) suggest a somewhat different circuit: specifically, they suggest stages of induction circuits, (1) assembling an encoding of each fewshot example’s input and moving it to fewshot label position, (2) attending to labels whose input is similar to the query, and copying these labels to the query. However, their experiments are restricted to semantic text classification (e.g., sentiment), in which fewshot inputs typically are either semantically aligned or opposed. While this is a reasonable circuit, the model would presumably use other circuit with function vectors for ICL tasks involving more diversified inputs. In addition, the arguments are largely based on evaluating representational similarity and on supervised probing, which does not establish that this mechanism is causally sufficient. While a causal intervention, through zeroing out the relevant components, shows this mechanism is relevant to ICL (Section 5.1 in Cho et al. (2025)), it leaves open if the mechanism alone can fully explain the behavior, or whether it is only a part of the overall mechanism.
Kharlapenko et al. (2025) perform a causal analysis leading to a sparse feature circuit in Gemma-2 2B. However, their circuit does not distinguish between inputs and outputs for different fewshot examples, making it hard to assess whether task information is created independently at each example or whether the representations are dependent across fewshot examples.
Overall, while function vectors play an important role for many ICL tasks, it remains unclear what information flow leads to aggregation of task information across fewshot examples. To foreshadow our results, we show that, while parallel aggregation of per-fewshot representations is important, the overall circuit is task-dependent, and may require a variety of additional components which, as we will show, contextualize the per-fewshot representations on other fewshot examples.
@@ 3 Results
We focus on Gemma-2, with a focus on its 2B version for tractability. This model had also been used by Kharlapenko et al. (2025), though our methodology allows us to also investigate larger variants (9B and 27B). We evaluated the model on tasks from Todd et al. (2024), and determined five tasks on which the 2B model achieves high (between $88\%$ and $100\%$) accuracy: Copying  [^1] This is not included in Todd et al. (2024), but we added it due to its naturalness., Present-Past, Capitalization, Country-Capital, Person-Sport (Figure 2). [^2] We also evaluated six other models at the 2B/3B scale, finding that they overall underperformed Gemma-2 2B at 10 shots on these tasks (Table 7), further motivating our focus on Gemma-2. We decided to focus on tasks with high accuracy in order to obtain a clear signal, as for these tasks there is a clear understanding of the input-output behavior exhibited by the full model. We focused experiments on 3-shot and 10-shot prompts, covering both shorter and longer prompts.
Each of the five tasks has an input space $\mathcal{X}$ (e.g., arbitrary lowercase words for Capitalization; countries for Country-Capital) and an output space $\mathcal{Y}$ (e.g., arbitrary capitalized words for Capitalization; cities for Country-Capital), and defines a function $f:\mathcal{X}\rightarrow\mathcal{Y}$ (e.g., mapping countries to their capitals) (Figure 25).
Figure 2: Accuracy of circuits (Gemma-2 2B, 9B, 27B). The Parallel circuit underperforms the full model by a large margin in some tasks. Contextualizing edges recover a large part of the gap. Whereas the Parallel already performs well in some tasks, we note that the contextualizing edges become crucial when evaluating on a difficult subset designed to be ambiguous between two tasks (Section 3.3).
@@ 3.1 Identifying Circuits
We investigate circuits on the level of the individual positions in a prompt. Whereas Kharlapenko et al. (2025) had collapsed different $x_{i}$’s into a single node (and same for all $y_{i}$’s), we keep them distinct, which allows us to understand information flow between fewshot examples.
We first discover a position-level circuit. Nodes are the positions in the prompt; edges are directed arcs between them, representing the flow of information (Figure 1). Position-level circuits provide useful and interpretable upper bounds on information flow, though they do not distinguish between computations happening in different layers. We treated all tokens within a fewshot input $x_{i}$ or output $y_{i}$ together, as the number of tokens varies between fewshot examples. As positions, we distinguish the fewshot inputs $x_{i}$ and outputs $y_{i}$ and the separators within ($t_{i}$) and between ($n_{i}$) examples; though $n_{i}$ play no role in any of the identified circuits.
@@
To identify circuits, we define a simple class of counterfactual inputs where all fewshot inputs and outputs (including the query $x_{N+1}$) are replaced with random words (Appendix G). The corrupted prompts still contain the same separators, and maintain the overall number of tokens of each $x_{i}$/$y_{i}$ (Figure 22). When we ablate an edge from position A to position B, the key (K) and value (V) activations of A when queried by B are replaced with activations computed on a corrupted prompt. The K/V activations of A when queried by other positions, as well as Q activation of B remain clean (Figure 28). This patching is applied simultaneously at each layer and head. We define the output of a circuit on an input prompt by greedily decoding a response $y_{N+1}$ until a separator is generated, while ablating all edges outside the circuit. [^3] We do, however, retain edges supporting the autoregressive prediction of the answer ($x_{N+1}\rightarrow y_{N+1}$ and $t_{N+1}\rightarrow y_{N+1}$) (Appendix D).
Parallel Circuit: Computing representations at each example and aggregate
@@
Edges between different fewshot examples
We next investigated which edges are needed beyond the Parallel subcircuit, aiming to find a small circuit recovering $\geq 90\%$ of the accuracy of the full model. We first ablated only edges involving non-final separators from the full model, and found this to do little harm across the five tasks compared to the full model, especially at 10 shots and for the 9B model (Tables 2, 5, 6), showing that relevant task information need not be routed through the separators in and between the fewshots, i.e., the separators may provide information about the presence of a prompt template, but do not play a nonredundant role in assembling information about the task. We thus continued with these edges ablated. We next considered the remaining logical possibilities, grouped by the types (input, output, separator) of positions: within-type edges (1) $x_{i}\rightarrow x_{j}$ (for all $1\leq i<j\leq N+1)$, (2) $y_{i}\rightarrow y_{j}$ (for all $1\leq i<j\leq N$), and across-type edges: (3) $x_{i}\rightarrow y_{j}$ ($1\leq i<j\leq N$), (4) $x_{i}\rightarrow t_{N+1}$ ($1\leq i\leq N$), (5) $y_{i}\rightarrow x_{j}$ ($11\leq i<j\leq N+$ 1). We investigated adding each of these groups individually to the aggregation circuit (Table 1). We found that adding (1) and (2) provided strong accuracy gains on some tasks, and focused on these two groups.
We next focused on different circuits involving these edges. We considered both general edge sets of types (1) and (2), and the local subset ($x_{i}\rightarrow x_{i+1}$ and $y_{i}\rightarrow y_{i+1}$); the latter were sufficient in all but one case ($y_{i}\rightarrow y_{j}$ needed for Capitalization). For each task, we chose the simplest circuit achieving $\geq 0.9$ at $N=3$, or (if there is none) the circuit achieving highest accuracy at $N=3,10$, all based on the 2B model’s accuracies (Tables 2). We illustrate the result for Country-Capital in Figure 1; circuits for all five tasks are shown in Figure 24. At 2B parameters and 10 fewshots, the circuits explained at least 90% of the full model’s performance, far exceeding the aggregation-only circuit on some tasks (Figure 2). We next evaluated the circuits on Gemma-2 9B and Gemma-2 27B, using the same methodology (Tables 5, 6). At 10 shots, the circuits that we had chosen based on the 2B model explained at least 90% of the full model’s performance in these model. Importantly, the Parallel circuit again underperformed the circuits involving contextualization on most tasks (Figure 2).
@@
In order to identify the heads responsible for the information flow we have identified, we next projected the position-level circuits to activation-level circuits. Here, nodes are activations of attention heads indexed by both their positions and layers; edges are directed arcs between the input and output of individual attention heads. More precisely, heads can be defined as tuples of (layer, head, position), and edges can be described as (layer, head, start position, end position). We used a gradient-based method (Michel et al., 2019; Syed et al., 2023) (see Appendix E for technical details) to obtain a circuit on the level of activations with edges defined by attention heads, in the 3-shot setup (Figures 913). The circuit reveals that the contextualization step tends to precede the aggregation step: contextualization happens in the lower half layers, most aggregation edges are in the middle and upper layers. Recall that prior work has established the existence of function vector heads, specific heads aggregating causally relevant task information at $t_{N+1}$ (Section 2). We next identified function vector heads by using the method described in Todd et al. (2024) for each task independently and selecting the top-10-scoring heads for each task. These heads were largely part of the identified circuits and mostly contributed to the $y_{i}\rightarrow t_{N+1}$ edges (Figure 26). This confirms that the circuits we have identified describe the buildup of task information up to the function vector heads.
Figure 3: Patching experiments manipulating the information about the input available to each edge in the full circuit, at the example of the Country-Capital task, 10-shot. See Figure 8 for results from all tasks.

### v3

@@ Abstract
In-Context Learning (ICL) is an intriguing ability of large language models (LLMs). Despite a substantial amount of work on its behavioral aspects and how it emerges in miniature setups, it remains unclear which mechanism assembles task information from the individual examples in a few-shot prompt. We use causal interventions to identify information flow in Gemma-2 2B for five naturalistic ICL tasks. We find that the model infers task information using a two-step strategy we call contextualize-then-aggregate: In the lower layers, the model builds up representations of individual few-shot examples, which are contextualized by preceding examples through connections between few-shot input and output tokens across the sequence. In the higher layers, these representations are aggregated to identify the task and prepare prediction of the next output. The importance of the contextualization step differs between tasks, and it may become more important in the presence of ambiguous examples. Overall, by providing rigorous causal analysis, our results shed light on the mechanisms through which ICL happens in language models.  [^1] https://github.com/lacoco-lab/icl_circuits
@@ 1 Introduction
In-Context Learning (ICL) is an intriguing property of large language models and has spurred a substantial amount of interest into how transformers are able to perform it (e.g. Brown et al., 2020; Min et al., 2022; Garg et al., 2022; Akyürek et al., 2023; Cho et al., 2025; Wang et al., 2023). Prior work in mechanistic interpretability has found function vectors (Todd et al., 2024; Hendel et al., 2023): attention heads whose output encodes task information, and which are causally responsible for the prediction of the response. However, the circuit by which these vectors are assembled remains only partly understood. Recent work has proposed that each few-shot example’s output token computes task information, which is then aggregated by attention heads to predict the next output (Wang et al., 2023; Cho et al., 2025; Kharlapenko et al., 2025), but this strategy has also been observed to leave part of the models’ performance unexplained (Cho et al., 2025).
In this paper, we use causal interventions (e.g. Vig et al., 2020; Geiger et al., 2021; Meng et al., 2022) to identify a circuit performing ICL on five naturalistic tasks formatted as few-shot prompts (Capitalization, Country-Capital, Present-Past, Person-Sport, Copying) in Gemma-2 2B (Team et al., 2024). We replicate the relevance of the aggregation step suggested by prior work, but show that it explains only a portion of the model’s full performance. We use causal interventions to identify information flow that recovers at least 90% of the model’s performance. A key idea is to first identify a computation graph between the tokens in a prompt (Figure 1). We find that the required graph differs between tasks: whereas the simplest circuit is largely sufficient on some tasks, others require additional computation paths. We find that the representations of few-shot examples later in the prompt are contextualized by information from prior few-shot examples, particularly, the last preceding few-shot example. Using causal interventions, we show that contextualization transports information about the input and output spaces, and the task itself. In particular,
  • Aggregation edges generally move information about the task from individual examples to the last token. In the Person-Sport task, they move information about input/output spaces without regard to functional relationship.
  • Contextualization edges generally move information about input/output spaces, and sometimes the task.
@@
    Obtain causally faithful information flow from the few-shot examples to the next-token prediction, explaining $\geq 90\%$ of the model’s performance.
    A key idea here is to focus on position-level circuits, where information flows only between a restricted subset of positions in a prompt, which is key to making the resulting computation graph interpretable.
@@
Figure 1: An NN-shot ICL prompt (here, $N=2$), at the example of the Country-Capital task. The green edges describe a Aggregation subcircuit that assembles information at each few-shot example in parallel ($x_{i}\rightarrow y_{i}$) and then aggregates these at $t_{N+1}$. The blue edges describe a Contextualization subcircuit that contextualizes the representation of each few-shot. A circuit involving both components recovers most of Gemma2-2B’s performance; aggregating without contextualization leads to a breakdown in performance on this task. See Figure 24 for other tasks.
@@ 2 Background: Aggregation and Function Vectors
Todd et al. (2024) (followed by Yin and Steinhardt (2025)) operationalize Function Vector Heads as heads whose activations, when patched from valid ICL prompts to shuffled uninformative prompts, can most increase the likelihood of the desired target (i.e., better performing the task articulated by the valid prompts). Aggregating activations from these attention heads results in function vectors, which can trigger execution of the task in uninformative or zero-shot contexts. Relatedly, Kharlapenko et al. (2025) find sparse autoencoder (SAE) features that encode task information (in few-shot examples) and features that execute them (at $t_{N+1}$). However, it remains largely open what circuit assembles this information, and how information is assembled from the different few-shot examples to these function vectors. One simple hypothesis could be that each few-shot example computes a representation, and this is aggregated by Function Vector Heads.
On the other hand, function vectors might not play an important role in certain kind of ICL tasks. Cho et al. (2025) suggest a somewhat different circuit: specifically, they suggest stages of induction circuits, (1) assembling an encoding of each few-shot example’s input and moving it to few-shot label position, (2) attending to labels whose input is similar to the query, and copying these labels to the query. However, their experiments are restricted to semantic text classification (e.g., sentiment), in which few-shot inputs typically are either semantically aligned or opposed. While this is a reasonable circuit, the model would presumably use other circuit with function vectors for ICL tasks involving more diversified inputs. In addition, the arguments are largely based on evaluating representational similarity and on supervised probing, which does not establish that this mechanism is causally sufficient. While a causal intervention, through zeroing out the relevant components, shows this mechanism is relevant to ICL (Section 5.1 in Cho et al. (2025)), it leaves open if the mechanism alone can fully explain the behavior, or whether it is only a part of the overall mechanism.
Kharlapenko et al. (2025) perform a causal analysis leading to a sparse feature circuit in Gemma-2 2B. However, their circuit does not distinguish between inputs and outputs for different few-shot examples, making it hard to assess whether task information is created independently at each example or whether the representations are dependent across few-shot examples.
Overall, while function vectors play an important role for many ICL tasks, it remains unclear what information flow leads to aggregation of task information across few-shot examples. To foreshadow our results, we show that, while parallel aggregation of per few-shot representations is important, the overall circuit is task-dependent, and may require a variety of additional components which, as we will show, contextualize the per few-shot representations on other few-shot examples.
@@ 3 Results
We focus on Gemma-2, with a focus on its 2B version for tractability. Our methodology allows us to also investigate larger variants (9B and 27B). We evaluated the model on tasks from Todd et al. (2024), and determined five tasks on which the 2B model achieves high (between $88\%$ and $100\%$) accuracy: Copying  [^2] This is not included in Todd et al. (2024), but we added it due to its naturalness., Present-Past, Capitalization, Country-Capital, Person-Sport (Figure 2). [^3] We also evaluated six other models at the 2B/3B scale, finding that they overall underperformed Gemma-2 2B at 10 shots on these tasks (Table 7), further motivating our focus on Gemma-2. We decided to focus on tasks with high accuracy in order to obtain a clear signal, as for these tasks there is a clear understanding of the input-output behavior exhibited by the full model. We focused experiments on 3-shot and 10-shot prompts, covering both shorter and longer prompts.
Each of the five tasks has an input space $\mathcal{X}$ (e.g., arbitrary lowercase words for Capitalization; countries for Country-Capital) and an output space $\mathcal{Y}$ (e.g., arbitrary capitalized words for Capitalization; cities for Country-Capital), and defines a function $f:\mathcal{X}\rightarrow\mathcal{Y}$ (e.g., mapping countries to their capitals) (Appendix, Figure 25). Each $N$-shot prompt then has the following structure: $x_{1}t_{1}y_{1}n_{1}\dots x_{N}t_{N}y_{N}n_{N}x_{N+1}t_{N+1}$, $\forall i:x_{i}\in\mathcal{X},y_{i}\in\mathcal{Y},y_{i}=f(x_{i}),t_{i}=\text{\textbackslash t},n_{i}=\text{\textbackslash n}$.
Figure 2: Accuracy of circuits (Gemma-2 2B, 9B, 27B). The Aggregation circuit underperforms the full model by a large margin in some tasks. Contextualizing edges recover a large part of the gap. Whereas the Aggregation already performs well in some tasks, we note that the contextualizing edges become crucial when evaluating on a difficult subset designed to be ambiguous between two tasks (Section 3.3).
@@ 3.1 Identifying Circuits
To localize the behavior of the model, we use patching, an approach widely adopted in the literature (Wang et al., 2022; Hanna et al., 2023). With this technique, a model is viewed as a computation graph with activations in different layers as nodes, and computations between them as edges. Then, it is possible to ablate some of the edges in the graph, forcefully replacing the computation along a specific edge with the one computed on the counterfactual input. Counterfactual inputs are designed to erase relevant information from the prompt while leaving other information intact for even better localization of model behavior. For example, when substituting "Berlin" with counterfactual "Paris", we isolate computations specific to the city’s identity, not those activated by its category (city) or token-type (word). If ablating a set of edges does not lead to a drop in the model’s performance, then the information unique to the original input relative to the counterfactual input, which was transferred along these edges, is not causal for the model’s prediction. This allows to discover the circuits - subgraphs of the full model’s computation graph that can explain a major part of model’s performance on a specific task.
In contrast to the standard approach, we focus on the information flow between positions in a prompt; consequently, we differentiate between activations of the same heads at different positions. We keep nodes for $x_{i}$ and $y_{i}$ in different examples distinct, which allows us to analyze the information flow between few-shot examples. [^4] This contrasts with the approach of Kharlapenko et al. (2025), who collapse all $x_{i}$ nodes into a single node (and same for all $y_{i}$ nodes).
We first discover a position-level circuit. Nodes are the positions in the prompt; edges are directed arcs between them, representing the flow of information (Figure 1). Importantly, in position-level circuits we collapse all edges across layers into a single edge. Ablating such an edge means ablating all corresponding edges in every attention head and layer. For illustration, $x_{2}$ in a position-level circuit can only receive four incoming edges: from $x_{1},t_{1},y_{1}$, or $n_{1}$.  [^5] The edges from/to bos/eos tokens are always ablated (Appendix E) Position-level circuits provide useful and interpretable upper bounds on information flow, though they do not distinguish between computations happening in different layers. We treated all tokens within a few-shot input $x_{i}$ or output $y_{i}$ together, as the number of tokens varies between few-shot examples. As positions, we distinguish the few-shot inputs $x_{i}$ and outputs $y_{i}$ and the separators within ($t_{i}$) and between ($n_{i}$) examples; though $n_{i}$ play no role in any of the identified circuits.
@@
To identify circuits, we define a simple class of counterfactual inputs where all few-shot inputs and outputs (including the query $x_{N+1}$) are replaced with random words (Appendix H). The corrupted prompts still contain the same separators, and maintain the overall number of tokens of each $x_{i}$$y_{i}$ (Appendix H, Figure 22). When we ablate an edge from position A to position B, the key (K) and value (V) activations of A when queried by B are replaced with activations computed on a corrupted prompt. The K/V activations of A when queried by other positions, as well as Q activation of B remain clean (Appendix H, Figure 28). This patching is applied simultaneously at each layer and head. We define the output of a circuit on an input prompt by greedily decoding a response $y_{N+1}$ until a separator is generated, while ablating all edges outside the circuit. [^6] We do, however, retain edges supporting the autoregressive prediction of the answer ($x_{N+1}\rightarrow y_{N+1}$ and $t_{N+1}\rightarrow y_{N+1}$) (Appendix E).
Aggregation Circuit: Computing representations at each example and aggregate
@@
Contextualization Subcircuit: Edges between different few-shot examples
We next investigated which edges are needed beyond the Aggregation subcircuit, aiming to find a small circuit recovering $\geq 90\%$ of the accuracy of the full model. We call all the edges in this circuit except those in the Aggregation subcircuit a Contextualization subcircuit. Specifically, the Contextualization subcircuit comprises edges that contextualize the representation of $(x_{i},y_{i})$ using prior few-shot examples $(x_{j},y_{j})$, where $j<i$. These edges connect nodes $A\rightarrow B$, with $A\in\{x_{j},y_{j}|j<i\}$ and $B\in\{x_{i},y_{i}\}$ for $i\leq N$. Note that the exact edges in this subcircuit may vary slightly across tasks.
We first ablated only edges involving non-final separators from the full model, and found this to do little harm across the five tasks compared to the full model, especially at 10 shots and for the 9B model (Tables 2, 5, 6), showing that relevant task information need not be routed through the separators in and between the few-shots, i.e., the separators may provide information about the presence of a prompt template, but do not play a nonredundant role in assembling information about the task. We thus continued with these edges ablated. We next considered the remaining logical possibilities, grouped by the types (input, output, separator) of positions: within-type edges (1) $x_{i}\rightarrow x_{j}$ (for all $1\leq i<j\leq N+1)$, (2) $y_{i}\rightarrow y_{j}$ (for all $1\leq i<j\leq N$), and across-type edges: (3) $x_{i}\rightarrow y_{j}$ ($1\leq i<j\leq N$), (4) $x_{i}\rightarrow t_{N+1}$ ($1\leq i\leq N$), (5) $y_{i}\rightarrow x_{j}$ ($1\leq i<j\leq N+1$). We investigated adding each of these groups individually to the aggregation circuit (Table 1). We found that adding (1) and (2) provided strong accuracy gains on some tasks, and focused on these two groups.
We next focused on different circuits involving these edges. We considered both general edge sets of types (1) and (2), and the local subset ($x_{i}\rightarrow x_{i+1}$ and $y_{i}\rightarrow y_{i+1}$); the latter were sufficient in all but one case ($y_{i}\rightarrow y_{j}$ needed for Capitalization). For each task, we chose the simplest circuit achieving $\geq 0.9$ at $N=3$, or (if there is none) the circuit achieving highest accuracy at $N=3,10$, all based on the 2B model’s accuracies (Tables 2). We illustrate the result for Country-Capital in Appendix, Figure 1; circuits for all five tasks are shown in Appendix, Figure 24. At 2B parameters and 10 few-shots, the circuits explained at least 90% of the full model’s performance, far exceeding the aggregation-only circuit on some tasks (Figure 2). We next evaluated the circuits on Gemma-2 9B and Gemma-2 27B, using the same methodology (Tables 5, 6). At 10 shots, the circuits that we had chosen based on the 2B model explained at least 90% of the full model’s performance in these model. Importantly, the Aggregation circuit again underperformed the circuits involving contextualization on most tasks (Figure 2).
@@
In order to identify the heads responsible for the information flow we have identified, we next projected the position-level circuits to activation-level circuits. Here, nodes are activations of attention heads indexed by both their positions and layers; edges are directed arcs between the input and output of individual attention heads. More precisely, heads can be defined as tuples of (layer, head, position), and edges can be described as (layer, head, start position, end position). We used a gradient-based method (Michel et al., 2019; Syed et al., 2023) (see Appendix F for technical details) to obtain a circuit on the level of activations with edges defined by attention heads, in the 3-shot setup (Figures 913). The circuit reveals that the contextualization step tends to precede the aggregation step: contextualization happens in the lower half layers, most aggregation edges are in the middle and upper layers. Recall that prior work has established the existence of function vector heads, specific heads aggregating causally relevant task information at $t_{N+1}$ (Section 2). We next identified function vector heads by using the method described in Todd et al. (2024) for each task independently and selecting the top-10-scoring heads for each task. These heads were largely part of the identified circuits and mostly contributed to the $y_{i}\rightarrow t_{N+1}$ edges (Appendix, Figure 26). This confirms that the circuits we have identified describe the buildup of task information up to the function vector heads.
Figure 3: Patching experiments manipulating the information about the input available to each edge in the full circuit, at the example of the Country-Capital task, 10-shot. For each ablation, we indicate whether the drop in performance is statistically significant compared to the full circuit performance according to a binomial test with $\alpha=0.05$. See Figure 8 for results from all tasks.